A body starting from rest acquires a velocity of 18 cm/s in 3 𝑠. If the body has a mass of 15 𝑔. What force is exerted on the body?


1
Expert's answer
2021-10-06T15:53:47-0400

Let's first find the acceleration of the body:


"a=\dfrac{v-v_0}{t}=\dfrac{18\ \dfrac{cm}{s}\cdot\dfrac{1\ m}{100\ cm}-0}{3\ s}=0.06\ \dfrac{m}{s^2}."

Finally, we can find the force exerted on the body from the Newton's Second Law of Motion:


"F=ma=0.015\ kg\cdot0.06\ \dfrac{m}{s^2}=9.0\cdot10^{-4}\ N."